OpenStudy (anonymous):

I have no clue what I'm doing wrong. I've followed a Master It example and a Watch It example given by webassign and I still come out with the same answer for: Find the average value f_ave of the function f on the given interval. f(x)=14x-x^4, [0,5] (work will be posted below)

OpenStudy (anonymous):

\[f_{ave}= \frac{ 1 }{ 5 }\int\limits_{0}^{5}(14x-x^2)dx= \frac{ 1 }{ 5 }[7x^2-\frac{ x^3 }{ 3 }]^5_0=\frac{ 1 }{ 5 }[175-\frac{ 125 }{ 3 }]=\frac{ 1 }{ 5 }[\frac{ 50 }{ 3 }]=\frac{ 10 }{ 3 }\]

OpenStudy (amistre64):

a - b/d is not equal to (a-b)/d

OpenStudy (anonymous):

That's how it was done on the example given by webassign but figures, half of the answers they give my professor change so it'll only take it in a different form.

OpenStudy (amistre64):

how do we subtract when dealing with fractions?

OpenStudy (anonymous):

To be honest, I have no clue. My professor just gives us the answers instead of working through the problems and showing my class how to do stuff.

OpenStudy (amistre64):

your dealing with calculus problems here .. not basic arithmetic. you should already be familiar with how to add/subtract fractions.

OpenStudy (anonymous):

Well I'm not.

OpenStudy (amistre64):

2 + 3/4 is not equal to 5/4

OpenStudy (anonymous):

Where the heck did you get 2+3/4 from?

OpenStudy (amistre64):

its an example ...

OpenStudy (anonymous):

OK...

OpenStudy (anonymous):

At least put it in terms of what the function is in?

OpenStudy (amistre64):

i did ...

OpenStudy (amistre64):

what are you doing wrong? your not working the fractions correctly

OpenStudy (anonymous):

Like I said above, I was going by examples that were given by webassign.

OpenStudy (amistre64):

what is: 175- 125/3

OpenStudy (anonymous):

175 is from 5 being put into the antiderivative of 14x which is 7x^2, thats where 175 came from and 125 is five put into the antiderivative of x^2 which is x^3/3, and that's where 125/3 came from.

OpenStudy (amistre64):

yes, and when we work the subtraction ... what is: 175 - 125/3 ???? it is not 50/3

OpenStudy (anonymous):

The examples given, heck even my professor, left the bottom 3 alone, that's where 50/3 came from.

OpenStudy (amistre64):

there is no 'bottom 3' common to both terms ....

OpenStudy (anonymous):

Well then my professor with 3 degrees in math is a fluttering idiot.

OpenStudy (amistre64):

i dont think thats the issue here.

OpenStudy (anonymous):

He teaches us with Theorems. He doesn't work the the problems, he doesn't tell us how to any of this. He just assumes that we know how to this stuff when in fact we do not.

OpenStudy (anonymous):

out*

OpenStudy (amistre64):

\[a+\frac bd\] \[a*\frac{d}{d}+\frac bd\] \[\frac{ad}{d}+\frac bd\] \[\frac{ad+b}{d}\ne\frac{a+b}{d}\]

OpenStudy (amistre64):

yes, he assumes you know how to operate on fractions seeing how you are dealing with a level of math that is far beyond the basics of arithmetic.

OpenStudy (amistre64):

\[175-\frac{125}{3}=\frac{3(175)-125}{3}\] \[3(175)-125\ne 50\]

OpenStudy (anonymous):

That still doesn't help me understand how to do this....